what is the simplified form of (3n)squared=360
Found 2 solutions by checkley71, troyapplehelen:
Answer by checkley71(8403)

3N~2=360 OR N~2=360/3 OR N~2=120 OR N=SQRT 120 OR N=2SQRT30

Answer by troyapplehelen(46)

1. 3n=a square root of 360 6x(square root of 10)
2. then n=6x(a square root of 10) over 3
=2(a square root of 10)